OpenStudy (anonymous):

The ddiagram drawn to scale shows the path by a boat travelling from A to B from B TO C and from C TO D. The trip takes 5h and 30minutes.

OpenStudy (anonymous):



OpenStudy (anonymous):

scale is 1cm =20km. a)Find the distance travelled by the boat b)find the magnitude of displacement of the boat c)determine the magnitude of the of the average velocity of the boat d)convert ans to m/s

OpenStudy (festinger):

a) Just measure the length of A to B and B to C (in cm). Add these 2 values together and multiply by 20 000 to get the distance in meters b) with the values of AB and BC, apply Pythagoras theorem to obtain the value for the hypotenuse. Multiply this value by 20 000 to get the value in meters. c) For magnitude, take the answer in (a) and divide by 19800s to get the answer in m/s. For average velocity, take b) and divide by 19800s. I am not sure if you are required to state the value for angle for velocity d) ok I think i messed up but you should be able to do the rest (at least get the idea) on how to solve questions like these. ask if you don't understand any of the steps.

OpenStudy (anonymous):

howd u get 19800

OpenStudy (anonymous):

@douglaswinslowcooper can u check this out 4 me ,to make sureimon the righttrack. my AB is 3.3 sm and my BC IS 11CM MY CD IS 2.8 when measured a) i got3.2+11+2.8=17 17x20=340 b)i gor 3^2+11^2= square root 131.24 =11.456 c)I am uncertain after this.:(

OpenStudy (anonymous):

a. should have three components, measured in cm then multiplied by 20km/cm. You seem to have done this. b) Displacement is just the distance from A to D, a straight line, same length as BC, and again multiply by 20km/cm c) average speed would be the total path/time, but average velocity is the displacement/time. You start with km and divide by the number of seconds, and then multiply this by 1000 m/km.

OpenStudy (anonymous):

so my a is right what about b?

OpenStudy (festinger):

I got 19800s by "counting" how many seconds there are in 5 hours and 30 minutes. 1 hour has 60 minutes, each minute has 60 seconds or, in math, 60*60=3600 the total time taken is 5.5 hours, so 5.5*3600 = 19800s